Movie tickets are $4 for regular admission and $3 for seniors. On a night where there were 190 people and total sales of $720, how many regular admission tickets were sold?
Answer by dragonwalker(73)

Let us say that the total number of regular admission tickets is a and the total number of senior tickets is s.
So each sale of one of these tickets brings in $4 and $3 respectively
So:
(a x 4) + (s x 3) is the total sales which is $720
i.e.
4a + 3s = 720
We also know the total number of people was 190 so:
a + s = 190
So if we rearrange the second equation we can get:
s = 190 - a
Now we substitute this into the first equation:
4a + 3(190 - a) = 720
4a + (570 - 3a) = 720
4a + 570 - 3a = 720
4a - 3a = 720 - 570
a = 150
so 150 regular admission tickets were sold !!!